Government Expenditures, G Tax Revenues, T Real GDP
180 100 500
180 120 600
180 140 700
180 160 800
180 180 900

Instructions: Enter your answers as whole numbers.


a. Waxwania is producing $600 of real GDP, whereas the potential real GDP (or full-employment real GDP) is $700. How large is its budget deficit?

$.

     How large is its cyclically adjusted budget deficit?
    
    $.

b. How large is its cyclically adjusted budget deficit as a percentage of potential real GDP?

Instructions: Round your answer to 2 decimal places.

   Deficit = percent.

c. Is Waxwania’s fiscal policy expansionary or is it contractionary?
Answer #1

Answer:

A] At $600 real GDP, the budget deficit =Government expenditure – Tax revenue
= $180-$120 = $60 billion
For potential real GDP of $700,
The cyclically-adjusted budget deficit = 180-140 = $40 billion


B] The cyclically-adjusted budget deficit as a percentage of potential real GDP = (40/700)*100 =5.714%


C] Since the government is running a cyclically-adjusted budget deficit, this fiscal policy is expansionary